OpenStudy (anonymous):

URGENT!! Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1. Answer choices: a ) f(x) = -1/4 x^2 b) f(x) = 1/4 x^2 c) f(x) = -4x^2 d) f(x) = 4x^2

OpenStudy (sosa4954):

I think it is focus at (2, −1) and a directrix of y = -1/2 This tells us the parabola opens downward as directrix is above focus also tells us the V(2, -3/4) %28-1-1%2F2%29%2F2+=+-3%2F4 where p = -1/4 a = 1/(4p), a = 1/(4(-1/4)) = -1

OpenStudy (anonymous):

so, the answer is a?

OpenStudy (sosa4954):

yea

OpenStudy (anonymous):

ok thank you

OpenStudy (sosa4954):

welcome